OpenStudy (vennestudent):

I need to find how many bricks they would need to build a walkway around a circular pool. The bricks are 4"x8" and the pool has a radius of 15 without the walkway. With the walkway it is 20. How many bricks would they need? I'm really lost

OpenStudy (sshayer):



OpenStudy (sshayer):

\[no.~of ~bricks=\frac{ 175\times 12 \times 12 }{ 32 }\times \frac{ 22 }{ 7 }=?\]